Last updated on April 15, 2017 at 6:47PM
ENCS server migrations for April 14th are completed. Services are operational.
ENCS server migrations for today are completed. Services are operational.
This is a reminder that ENCS and Datacenter staff will begin migrating equipment and services to Datacenter facilities beginning on Monday April 3rd, and continuing through April 14th. Work will be beginning at 4:00 AM and concluding by 7:00 AM each day.
Again, physical moves are planned on Monday April 3rd and Wednesday April 5th. During the hours mentioned above, there may be brief interruptions in the following:
otc2.psu.edu (22.214.171.124) -- Caching DNS Server, clients with multiple DNS servers will continue to function, but at a slower rate.
clock.psu.edu — NTP clients will continue to function normally.
Beginning Monday April 3rd and continuing through April 14th ENCS and Datacenters staff relocate equipment and services. Work will begin at 4:00 AM and conclude by 7:00 AM each day. Measures are being taken to minimize the impact of affected services during this timeframe. This work is being tracked under CHG0040544.
The physical moves are planned on Monday April 3rd and Wednesday April 5th. During the hours mentioned above, there may be brief interruptions in the following:
The transfer of additional services to VMHosting are scheduled from Aril 6th — April 14th -- no significant outages are expected during these moves.
For more information, please contact Operations Center (814-865-4662).